OpenStudy (firejay5):

Find both the complement and supplement of each angle measure. Show and explain your work. 24. y degrees 26. (x + 30) degrees 28. (90 - z degrees) My Answers: I feel some are wrong but who knows. 24. 90 degrees; 0 degrees 180 degrees; 0 degrees 26. x + 30 = 90 x = 60 degrees x + 30 = 180 x = 150 degrees 28. (90 - z degrees) = 90 z = 0 degrees (90 - z degrees) = 180 -z = 90 z = -90 degrees

OpenStudy (dumbcow):

2 complimentary angles add up to 90 So angle y and its complement add up to 90 ----> y + a = 90 ----> a = 90 - y 90 -y would then by the compliment to angle y.

OpenStudy (firejay5):

Why wouldn't it be 90 degrees and 0 degrees, because it only gave you 1 angle y = 90 degrees, so than the other angle would be 0 degrees? @dumbcow

OpenStudy (dumbcow):

oh does it say that y=90 ? however, if we do not know y, then its just 90-y

OpenStudy (firejay5):

what about the other problems

OpenStudy (firejay5):

so it would be 90 - y and 180 - y

OpenStudy (dumbcow):

follow the same approach, instead of equaling it to 90. Subtract them from 90. Given angle a, find compliment c a+c = 90 c = 90 - a If a = x+30 Then, c = 90 - (x+30) = 60 - x

OpenStudy (dumbcow):

correct above

OpenStudy (firejay5):

24. 90 - y & 180 - y 26. 60 - x & 150 - x 28. 0 - z & -90 - z @dumbcow Cause it asks for the complement and supplement not just the complement

OpenStudy (dumbcow):

yep i was leaving the supplement for you since the method is same, just replace 90 with 180 24, 26 correct 28) you need check your neg signs z z+90

OpenStudy (firejay5):

okay so just leave the negative out

OpenStudy (dumbcow):

90 - (90-z) = 90-90-(-z) = z

OpenStudy (firejay5):

okay what???

OpenStudy (dumbcow):

when you distribute the neg, you get neg times neg, which equals a positive

OpenStudy (firejay5):

so it's 90 + z or -90 + z

OpenStudy (dumbcow):

180 - (90 - z) = 180 - 90 -(-z) = 90 + z

OpenStudy (firejay5):

thank ya so much for explaining it, it makes more sense to me now